A neutron star is the remnant left after certain supernovae (explosions of giant stars). Typically, neutron stars are about 18.0 km in diameter and have around the same mass as our sun. What is a typical neutron star density in  g/cm3?
Answer #1

given,

diameter of star = 18 km or 18 * 10^6 cm

mass = 1.989 × 10^30 kg or 1.989 × 10^33 gm

density = mass / volume

volume = (4/3) * pi * radius^3

volume = (4/3) * pi * (9 * 10^6)^3

density = 1.989 × 10^33 / ((4/3) * pi * (9 * 10^6)^3)

density = 651356341172 gm/cm^3